OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4 Choose the equation of the vertical line that passes through the point (4, -3). y = 4 x = - 3 y = -3 x = 4

OpenStudy (anonymous):

x = - 3

OpenStudy (anonymous):

and the other ?

OpenStudy (cwrw238):



OpenStudy (cwrw238):

the line passes through x = 4 and infinite values of y....

OpenStudy (anonymous):

so which answer choice is it ??

OpenStudy (cwrw238):

well x has a value and y can be anything so should we have x or y in the equation?

OpenStudy (anonymous):

i have no clue

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

its x = 4

OpenStudy (cwrw238):

everywhere along that line x = 4 y isn't included because it can be any value

OpenStudy (anonymous):

okay i see now but i thought we were doing the first one. im soo confused. im off for an hour plese explain the first one to me so i can give u best responce when i get back on.

OpenStudy (cwrw238):

the slope of the first line = -1/2 sand passes through (4,-6) use the general form y-y1 = m(x-x1) m = -1/2, x1= 4 and y1 = -6 so plug in these values y -(-6) = =(-1/2)(x - 4) = (-1/2)x + 2 y = (-1/2)x + 8